OpenStudy (anonymous):

The Goodyear blimps, which frequently fly over sporting events, hold approximately 1.90×105ft3 of helium. If the gas is at 25∘C and 1.0 atm, what mass of helium is in the blimp? I really only need to know how to get the ft3 to liters please

OpenStudy (jfraser):

since\[1 ft = 30.48cm\]then\[(1ft)^3 = (30.48cm)^3\]but you have to cube everything, the number AND the unit, so \[1ft^3 = (30.48)^3cm^3\]or\[1ft^3 = 28316cm^3\]Also,\[1L = 1000cm^3\]

OpenStudy (anonymous):

Awesome, thank you :)